OpenStudy (anonymous):

Alex designed the model of a cylindrical water tank to be installed in a park. The model has a radius of 5 cm and a height of 14 cm. The dimensions of the actual water tanks that will be installed in the park are 8 times larger than the model. The total surface area of the actual water tanks, rounded to the nearest whole number, is _________ cm². Use 22/7 for pi.

OpenStudy (anonymous):

I KNOW I HAVE TO USE THE TOTAL SURFACE AREA FORMULA, BUT DONT KNOW WHICH ONE FOR THIS CYLINDRICAL SHAPE TANK @Calcmathlete

OpenStudy (anonymous):

SA = 2πr^2 + 2πrh

OpenStudy (anonymous):

Then use the ratio. If the dimensions will be 8 times larger, the ratio will be 1/64

OpenStudy (anonymous):

so new formula is SA= 2*22/7*5^2 + 2*22/7*5*14?

OpenStudy (anonymous):

Yes. Then you have to get the surface area of the actual tank using the ratio, 1/64.

OpenStudy (anonymous):

ok give me a minute ;)

OpenStudy (anonymous):

without doing the ratio part i got 597.142...the ratio part is the one I forgot how to do ;/

OpenStudy (anonymous):

can you explain how i can do the ratio part?

OpenStudy (anonymous):

Sure. It says that the dimensions are 8 times the ones on the model. However, the ratio of the areas is that SQUARED. Dimensions Ratio: a : b Area Ratio : a^2 : b^2 Therefore, set up a proportions. 1/64 = 597.142/x x = 597.142(64

OpenStudy (anonymous):

ok so is it 9.33? @Calcmathlete

OpenStudy (anonymous):

No. Multiply 597.142 and 64 together.

OpenStudy (anonymous):

38217